OpenStudy (anonymous):

Write an equation for the translation of x^2+y^2=25 by 7 units left and 2 units down. (x+7)^2+(y+2)^2=25 (x-7)^2+(y+2)^2=25 (x+7)^2+(y-2)^2=25 (x-7)^2+(y-2)^2=25

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Okayy, so we have the equation of \(\Large x^2+y^2=25\) correct? Do you know what the center of this equation is?

OpenStudy (anonymous):

No I don't. Will you help me?

OpenStudy (anonymous):

Actually is the center just (0,0)? Because that is the center of the equation.

OpenStudy (luigi0210):

In order to under it, you'll need to know the basics of the conic shapes. And yes, the center would be (0,0) with a radius of 5. Now, we want to move it 7 units to the left, and 2 units down, like so:  (pretend that circle is the center) When we do that, we are moving in the x-direction (positively) and y-direction (negatively).. which in a coordinate point is (7, -3) The radius won't change though So change the (0, 0) to (7, -3) in \(\Large (x-0)^2+(y-0)^2=25 \)

OpenStudy (luigi0210):

*(7, -2) not (7, -3) my mistake

OpenStudy (anonymous):

So C would be correct?

OpenStudy (luigi0210):

In short, \(\Large (\color{red}{7}, \color{blue}{-2}) \) will be the new \(\Large (\color{red}{h}, \color{blue}{k})\) And nope, try again: \(\Large (x-\color{red}{h})^2 +(y-\color{blue}{k}) =25 \) \(\Large (x-\color{red}{7})^2 +(y-\color{blue}{(-2)}) =25 \) Do understand it some at least?

OpenStudy (luigi0210):

Forgot the squares, but you get the point right?

OpenStudy (anonymous):

So the opposite. -7 +2?

OpenStudy (luigi0210):

Yup! Gotta watch out for those negatives.

OpenStudy (luigi0210):

It'll end up looking like this: http://prntscr.com/9o39l6